# Date of publication xxxx 00, 0000, date of current version xxxx 00, 0000. xx

### v2

@@ Abstract
This paper presents an infrastructure-free approach for obstacle detection and environmental mapping using ultra-wideband (UWB) radar mounted on a mobile robotic platform. Traditional sensing modalities such as visual cameras and Light Detection and Ranging (LiDAR) fail in environments with poor visibility due to darkness, smoke, or reflective surfaces. In these visioned-impaired conditions, UWB radar offers a promising alternative. To this end, this work explores the suitability of robot-mounted UWB radar for environmental mapping in dynamic, anchor-free scenarios. The study investigates how different materials (metal, concrete and plywood) and UWB radio channels (5 and 9) influence the Channel Impulse Response (CIR). Furthermore, a processing pipeline is proposed to achieve reliable mapping of detected obstacles, consisting of 3 steps: (i) target identification (based on CIR peak detection), (ii) filtering (based on peak properties, signal-to-noise score, and phase-difference of arrival), and (iii) clustering (based on distance estimation and angle-of-arrival estimation). The proposed approach successfully reduces noise and multipath effects, resulting in an obstacle detection precision of at least $90.71\%$ and a recall of $88.40\%$ on channel 9 even when detecting low-reflective materials such as concrete. This work offers a foundation for further development of UWB-based localisation and mapping (SLAM) systems that do not rely on visual features and, unlike conventional UWB localisation systems, do not require on fixed anchor nodes for triangulation.
@@ I Introduction
The ability to detect objects and build a representation of the surrounding environment is an essential requirement to allow autonomous operations of mobile robots. Existing systems rely on cameras to capture images and track movement by extracting visual features. However, this method depends on clear lighting conditions and struggles in environments with low visibility, shadows, or reflective surfaces [review-LiDAR]. Light Detection and Ranging (LiDAR) systems, on the other hand, use laser pulses to measure distances and construct maps. While highly accurate, LiDAR sensors face challenges in environments where laser beams are scattered or absorbed, such as in heavy rain, fog, or areas with multiple reflective surfaces. Given these challenges, there is a growing need for alternative localisation and mapping technologies. Ultra-wideband (UWB) is a promising alternative due to its ability to operate in low-visibility environments while maintaining high accuracy (typically under 20 cm). However, UWB-based localisation systems usually rely on anchor-tag setups, where fixed UWB anchors are installed in predefined locations, and mobile robots are equipped with UWB-tags [ensemble-learning] [nlos-neural-network]. Although this setup works well for environments with stable infrastructure, it is impractical in dynamic or unknown places, such as search and rescue activities or industrial facilities where the surroundings frequently change. Furthermore, using a UWB radio instead of LiDAR can reduce the system’s total weight, a significant advantage for drones that need to be as lightweight as possible to maximize autonomy [survey-robot-swarms].
To realize this vision, this paper proposes and evaluates signal processing methods to detect obstacles and map the environment. For evaluating the proposed algorithm, data was collected using the Qorvo QM33120WDK1 development kit [datasheet-sensors]. The setup consists of separate omnidirectional transmitter antenna and directional receiver antennas mounted on a TurtleBot 4. The main contributions of this paper are as follows:
@@
  • We broaden the radar’s field of view by clustering detected points using combined distance and angle-of-arrival (AoA) estimations, enabling spatially coherent obstacle mapping.
@@
The remainder of this paper is organised as follows: Section II provides background information on UWB. Section III discusses related work, whereas Section IV presents the proposed processing pipeline. In Section V, the data collection is explained, and the results for the proposed method are presented in Section VI together with a visual demonstration. Finally, Sections VII and VIII cover suggestions for future work and the conclusion, respectively.
@@ II Background of Ultra-Wideband
UWB is a radio technology that transmits data over a wide frequency range, either with an absolute bandwidth greater than 500 MHz or a fractional bandwidth ($B_{F}$) greater than 20%. The fractional bandwidth ($B_{F}$) is defined as the ratio of the signal’s bandwidth ($B_{W}$) to its centre frequency ($f_{C}$). The fractional bandwidth ($B_{F}$) is defined as:
@@
Where $f_{H}$ and $f_{L}$ are the upper and lower frequencies of the -10 dB bandwidth [uwb-overview].
@@
This technology operates within the frequency spectrum of 3.1 GHz to 10.6 GHz, enabling data transmission speeds of up to 110 Mbps, typically with a range of up to 10 meters. A higher range of up to 100 meters is also possible, but at lower data rates, depending on several factors, including channel frequency, antenna design, power levels, and complexity of the propagation environment [uwb-standards].
UWB radar can be implemented either as Direct Sequence UWB (DS-UWB), which spreads a carrier signal using pseudo-random noise code, or as Impulse Radio UWB (IR-UWB), which transmits short pulses and estimates distance using the Time-of-Flight (ToF). Compared to DS-UWB, IR-UWB is simpler to implement, more power-efficient, and more robust against multipath interference, which is why it is most used in radar applications [uwb-overview]. For this reason, this proposed approach focuses on pulse-based UWB.
To better understand how IR-UWB radar performs in realistic environments, it is essential to study how the transmitted signal interacts with obstacles and reflections in the environment. This is described by the Channel Impulse Response (CIR), illustrated in Fig. 1. The first peak in the figure corresponds to the Line-of-Sight (LoS) component, which is the earliest signal arriving directly from the transmitter to the receiver assuming no obstacles in between. In this case, the LoS signal indicates the signal transmitted from the transmitter antenna to the receiver antenna co-located on the same robot. The antennas are separated by a distance of 21 cm. Subsequent peaks represent reflections from other objects, walls, or multipath components.
@@
The CIR is typically modelled as a sum of deterministic and diffuse multipath components [cir-formula] as shown in (1):
@@
The first term represents the deterministic multipath components, which describe the propagation of the UWB signal through multipaths with different delays. Each component is characterized by an amplitude $A_{i}$ and a corresponding delay $\tau_{i}$. The second term accounts for the diffuse multipath components, capturing random variations and scattering effects within the channel. These are modelled as additive white Gaussian noise (AWGN), denoted by $\nu(t)$, to provide a realistic representation of the channel behaviour.
@@
UWB radar is widely used in applications that require precise distance measurements, even in cluttered or dynamic environments. Some examples of these applications are presence detection, device-free localisation, monitoring vital signs without physical contact [lambrecht2025lowcostembeddedbreathingrate] and, activity recognition. UWB is a strong candidate for fast-moving robots and dynamic environments due to its low weight, low power consumption, and cost-effectiveness [survey-robot-swarms]. While RGB cameras are also lightweight and energy-efficient, they suffer in complex or low-light environments. LiDAR offers high-accuracy long-range sensing but is heavy and power-intensive. LiDAR performance degrades in environments with smoke, dust, or fire, as laser beams cannot penetrate these particles. Additionally, both LiDAR and depth cameras have difficulties detecting transparent objects. Vanhie-Van Gerwen et al. [indoor-drone] show that across various drone sizes, UWB stands out for its favourable accuracy-to-cost ratio. It also provides precise positioning thanks to its wide bandwidth and short pulses, is resistant to narrowband interference, and respects privacy by not capturing identity-related data. Both mmWave and UWB can distinguish small objects thanks to their high bandwidth. However, signals with lower centre frequencies generally propagate farther. mmWave operates between 30 GHz and 300 GHz, resulting in high-speed data transmission, but the technology suffers from high path loss, limited penetration, and short range. In contrast, UWB experiences less attenuation over distance and better penetration of walls and obstacles.
@@ III Related Work
[cir-processing] Object/person detection using CIR analysis Three CIR processing techniques (ICIR, UCIR, ACIR) with BMA and BSDA algorithms DW1000 - - whiteboard, large tv, metal box, person Avg. error $\leq$ 9 cm using ACIR (larger for persons or smaller objects)
[machine-learning-train] Detect humans/obstacles in automatic train pairing. Supervised Machine Learning P440 from TimeDomain - - - - humans/obstacles Best detection accuracy: 55%-70% indoors,>> 95% outdoors
[distance-estimation-ir-uwb] Remove clutter to improve detection accuracy Low-rank approximation and singular value decomposition methods not specified - - - walking person Avg. error <20<20 cm, RMSE <50<50 cm
[LSTM-based-system-multiple-obstacle-detection] Distinguish objects from noise in transportation use cases. Uses Discrete Wavelet Transform to extract features and forward these to the LSTM network. UMAIN HST-D3 directional antenna - - - - pedestrian, cyclist, vehicle, tram 72.78% precision, 71.34% recall, 72.06% F1-score
[entropy-based] Detect the target object Utilises signal’s entropy UMAIN HST-D3 directional antenna - - - - pedestrian, cyclist, vehicle, tram 75.34% precision, 63.06% recall, 68.65% F1-score
[localisation-navigation-uwb] Signal processing chain for obstacle detection Background subtraction, CFAR and Kalman filter M-sequence UWB - - wall and glass pane not specified
[cooperative-uwb-slam] UWB-SLAM Kalman Filter to reduce followed by particle filter M-Sequence UWB - - walls and corners $\leq$ 20 cm
[anchorless-uwb-1] UWB-SLAM wheel odometry and Extended Kalman Filter X4M300 - - vertical metal rods 6.2 cm localisation error
[anchorless-uwb-2] UWB-SLAM in feature-deficient areas Extends previous approach with UWB AoA anchor-tag for better loop closure X4M300 - - - vertical metal rods 10.3 cm localisation error
This paper Reduce noise and unwanted multipath for reliable obstacle mapping Filtering on peak properties, SNR-score, PDoA and clustering QM33120WDK1 metal plate, concrete and plywood box MAE $\leq$ 11.50 cm, $\geq 90.71\%$ precision, $\geq 88.40\%$ recall on channel 9
@@
Table I shows a comparison of the existing work with our approach. Most state-of-the-art solutions rely on the installation of fixed UWB radar systems. For example, Van Herbruggen et al. [cir-processing] introduce Accumulation Channel Impulse Response (ACIR) technique, which improves signal resolution for obstacle detection. Although this method increases the detection accuracy, it requires the collection of sufficient samples before it can be applied, limiting the maximum speed of robotic platforms. From the research of Sattiraju et al. [machine-learning-train] it is concluded that Random Forest and Extra Trees have the highest accuracy in obstacle detection. However, these methods are prone to overfitting the training data, which makes the model generalise less in new environments. Yun et al. [distance-estimation-ir-uwb] suppress noise from the received signal using singular value decomposition (SVD). Although their method achieves high accuracy, the required matrix computations are computationally intensive and may increase the latency and memory usage for real-time mobile robots. Mimouna et al. [LSTM-based-system-multiple-obstacle-detection], [entropy-based] focus on obstacle detection, but lack characterisation of the obstacle (e.g. estimation of distance and width). Moreover, the use of complex deep learning models in [LSTM-based-system-multiple-obstacle-detection] would lead to a higher execution time. C. Smeenk et al. [localisation-navigation-uwb] present a signal processing chain for M-sequence UWB radars mounted on a mobile robot. Their approach combines background subtraction to remove the antenna crosstalk, with a constant false alarm rate (CFAR) method for object detection and distance estimation. However, the localisation is limited to reporting that an object is in front of the antenna and its distance, without providing its relative position in the environment. The method presented in this work overcomes all the above limitations by using clustering that is more adaptable to unseen and dynamic environments. Moreover, our work also characterizes the position and width of detected obstacles by combining both distance estimation and AoA information.
To the best of our knowledge, only four prior publications focus on robot-mounted UWB radar systems.
@@
[anchorless-uwb-1] and [anchorless-uwb-2] both focus more on Simultaneous Localisation and Mapping (SLAM) localisation aspects than UWB signal processing optimisations for obstacle characterisation. They are also limited by recognizing only vertical rods rather than reconstructing full obstacle representations from various material types. The proposed approach from our work can serve as a complementary method to [anchorless-uwb-1] and [anchorless-uwb-2], aiming to suppress noise and mitigate ghost detections while increasing the overall performance of their proposed UWB-SLAM algorithm. Unlike [anchorless-uwb-2], which complements UWB-SLAM with localisation by employing AoA nodes for robot localisation and loop closure, our paper focuses on estimating the AoA of detected objects using the PDoA between antennas on the same chip. R. Zetik et al. [cooperative-uwb-slam] also focus on SLAM and demonstrate good localisation accuracy across different room shapes. However, their approach relies on proprietary antenna designs tailored for their experiments, which increases deployment cost.
@@
Finally, almost all prior work relies on custom UWB radar-only hardware devices with larger bandwidths and thus higher range resolution. Instead, our proposed approach is fully compliant with low-cost IEEE 802.15.4 standard off-the-shelf hardware. Using this setup, the same hardware can also localise the robot if fixed infrastructure is available, by reusing the transmitted wireless signal for multiple purposes simultaneously. This results in reduced energy consumption, less need for separate hardware, lower costs and weight, and more efficient use of the sparse spectrum. In this way, low-cost joint sensing, communication, and localisation for robotic platforms can be realised [integrated-sensing-and-communication].
@@ IV Proposed approach
Figure 2: Overview of the proposed approach: Each capture contains the preamble CIR and two additional CIRs from both antennas on the same chip used for calculating the PDoA. A new measurement happens every 10.42 ms and contains the data of the environment at one timestamp. These are processed independently followed by filtering out noise or unwanted multipath components using peak properties, SNR-score, and PDoA. The distance to the object is estimated and the AoA is then calculated. Subsequently, multiple rows are combined to cluster reflections from the same object and map the environment.

### v3

@@ Abstract
This paper presents an infrastructure-free approach for obstacle detection and environmental mapping using ultra-wideband (UWB) radar mounted on a mobile robotic platform. Traditional sensing modalities such as visual cameras and Light Detection and Ranging (LiDAR) fail in environments with poor visibility due to darkness, smoke, or reflective surfaces. In these vision-impaired conditions, UWB radar offers a promising alternative. To this end, this work explores the suitability of robot-mounted UWB radar for environmental mapping in anchor-free, unknown scenarios. The study investigates how different materials (metal, concrete and plywood) and UWB radio channels (5 and 9) influence the Channel Impulse Response (CIR). Furthermore, a processing pipeline is proposed to achieve reliable mapping of detected obstacles, consisting of 3 steps: (i) target identification (based on CIR peak detection), (ii) filtering (based on peak properties, signal-to-noise score, and phase-difference of arrival), and (iii) clustering (based on distance estimation and angle-of-arrival estimation). The proposed approach successfully reduces noise and multipath effects, achieving high obstacle detection performance across a range of materials. Even in challenging low-reflectivity scenarios such as concrete, the method achieves a precision of $73.42\%$ and a recall of $83.38\%$ on channel 9. This work offers a foundation for further development of UWB-based localisation and mapping (SLAM) systems that do not rely on visual features and, unlike conventional UWB localisation systems, do not require fixed anchor nodes for triangulation.
@@ I Introduction
The ability to detect objects and build a representation of the surrounding environment is an essential requirement to allow autonomous operations of mobile robots. Existing systems rely on cameras to capture images and track movement by extracting visual features. However, this method depends on clear lighting conditions and struggles in environments with low visibility, shadows, or reflective surfaces [1]. Light Detection and Ranging (LiDAR) systems, on the other hand, use laser pulses to measure distances and construct maps. While highly accurate, LiDAR sensors face challenges in environments where laser beams are scattered or absorbed, such as in heavy rain, fog, or areas with multiple reflective surfaces. Given these challenges, there is a growing need for alternative localisation and mapping technologies. Ultra-wideband (UWB) is a promising alternative due to its ability to operate in low-visibility environments while maintaining high accuracy (typically under 20 cm). However, UWB-based localisation systems usually rely on anchor-tag setups, where fixed UWB anchors are installed in predefined locations, and mobile robots are equipped with UWB-tags [2] [3]. Although this setup works well for environments with stable infrastructure, it is impractical in unknown places, such as search and rescue activities or industrial facilities where the surroundings frequently change. Furthermore, using a UWB radio instead of LiDAR can reduce the system’s total weight, a significant advantage for drones to maximize their autonomy [4].
To realize this vision, this paper proposes and evaluates signal processing methods to detect obstacles and map the environment. For evaluating the proposed algorithm, data were collected using the Qorvo QM33120WDK1 development kit [5]. The setup consists of separate omnidirectional transmitter antenna and directional receiver antennas mounted on a TurtleBot 4. The main contributions of this paper are as follows:
@@
  • We improve UWB obstacle detection by combining angle-of-arrival (AoA) and distance to cluster radar returns consistently across time, producing stable and spatially coherent obstacle maps.
@@
The remainder of this paper is organised as follows: Section II provides background information on UWB. Section III discusses related work, whereas Section IV presents the proposed processing pipeline. In Section V, the data collection is explained, and the results for the proposed method are presented in Section VI together with a visual demonstration. Finally, Sections VII and VIII cover limitations and the conclusion with suggestions for future work, respectively.
@@ II Background of Ultra-Wideband
UWB is a radio technology that transmits data over a wide frequency range, either with an absolute bandwidth greater than 500 MHz or a fractional bandwidth ($B_{F}$) greater than 20%. The fractional bandwidth ($B_{F}$) is defined as the ratio of the signal’s bandwidth ($B_{W}$) to its centre frequency ($f_{C}$) as:
@@
Where $f_{H}$ and $f_{L}$ are the upper and lower frequencies of the -10 dB bandwidth [6].
@@
This technology operates within the frequency spectrum of 3.1 GHz to 10.6 GHz, enabling data transmission speeds of up to 110 Mbps, typically with a range of up to 10 meters. A higher range of up to 100 meters is also possible, but at lower data rates, depending on several factors, including channel frequency, antenna design, power levels, and complexity of the propagation environment [7].
UWB radar can be implemented either as Direct Sequence UWB (DS-UWB), which spreads a carrier signal using pseudo-random noise code, or as Impulse Radio UWB (IR-UWB), which transmits short pulses and estimates distance using the Time-of-Flight (ToF). Although pseudo-random noise radars are common in scientific research, they are typically not compatible with IEEE 802.15.4 UWB localisation and communication systems, and no low-cost commercial radio chips exist. For this reason, this proposed approach focuses on pulse-based UWB.
To better understand how IR-UWB radar works, it is essential to study how the transmitted signal interacts with obstacles and reflections in the environment. This is described by the CIR, illustrated in Fig. 1. The first peak in the figure corresponds to the Line-of-Sight (LoS) component. In this case, the LoS signal indicates the signal transmitted from the transmitter antenna to the receiver antenna co-located on the same robot. Subsequent peaks represent reflections from other objects, walls, or multipath components.
@@
The CIR is typically modelled as a sum of deterministic and diffuse multipath components [8] as shown in (1):
@@
The first term represents the deterministic multipath components, which describe the propagation of the UWB signal through multipaths with different delays. Each component is characterized by an amplitude $A_{i}$ and a corresponding delay $\tau_{i}$. The second term accounts for measurement noise, modelled as additive white Gaussian noise (AWGN), denoted by $\nu(t)$.
@@
UWB radar is widely used in applications that require precise distance measurements, even in cluttered or unknown environments. Some examples of these applications are presence detection, device-free localisation, monitoring vital signs without physical contact [9] and, activity recognition. UWB is a strong candidate for fast-moving robots and unknown environments due to its low weight, low power consumption, and cost-effectiveness [4]. While RGB cameras are also lightweight and energy-efficient, they suffer in complex or low-light environments. LiDAR offers high-accuracy long-range sensing but is heavy and power-intensive. LiDAR performance degrades in environments with smoke, dust, or fire, as laser beams cannot penetrate these particles. Additionally, both LiDAR and depth cameras have difficulties detecting transparent objects. Vanhie-Van Gerwen et al. [10] show that across various drone sizes, UWB stands out for its favourable accuracy-to-cost ratio. It also provides precise positioning thanks to its wide bandwidth and short pulses, is resistant to narrowband interference, and respects privacy by not capturing identity-related data. Both mmWave and UWB can distinguish small objects thanks to their high bandwidth. However, signals with lower centre frequencies generally propagate farther. mmWave operates between 30 GHz and 300 GHz, resulting in high-speed data transmission, but the technology suffers from high path loss, limited penetration, and short range. In contrast, UWB experiences less attenuation over distance and better penetration of walls and obstacles.
@@ III Related Work
[11] Object/person detection using CIR analysis Three CIR processing techniques (ICIR, UCIR, ACIR) with BMA and BSDA algorithms DW1000 - - whiteboard, large tv, metal box, person Avg. error $\leq$ 9 cm using ACIR (larger for persons or smaller objects)
[12] Detect humans/obstacles in automatic train pairing. Supervised Machine Learning P440 from TimeDomain - - - - humans/obstacles Best detection accuracy: 55%-70% indoors,>> 95% outdoors
[13] Remove clutter to improve detection accuracy Low-rank approximation and singular value decomposition methods not specified - - - walking person Avg. error <20<20 cm, RMSE <50<50 cm
[14] Distinguish objects from noise in transportation use cases. Uses Discrete Wavelet Transform to extract features and forward these to the LSTM network. UMAIN HST-D3 directional antenna - - - - pedestrian, cyclist, vehicle, tram 72.78% precision, 71.34% recall, 72.06% F1-score
[15] Detect the target object Utilises signal’s entropy UMAIN HST-D3 directional antenna - - - - pedestrian, cyclist, vehicle, tram 75.34% precision, 63.06% recall, 68.65% F1-score
[16] Signal processing chain for obstacle detection Background subtraction, CFAR and Kalman filter M-sequence UWB - - wall and glass pane not specified
[17] UWB-SLAM Kalman Filter to reduce followed by particle filter M-Sequence UWB - - walls and corners $\leq$ 20 cm
[18] UWB-SLAM wheel odometry and Extended Kalman Filter X4M300 - - vertical metal rods 6.2 cm localisation error
[19] UWB-SLAM in feature-deficient areas Extends previous approach with UWB AoA anchor-tag for better loop closure X4M300 - - - vertical metal rods 10.3 cm localisation error
This paper Reduce noise and unwanted multipath for reliable obstacle mapping Filtering on peak properties, SNR-score, PDoA and clustering QM33120WDK1 metal plate, concrete and plywood box MAE $\leq$ 16.70 cm, $\geq 73.42\%$ precision, $\geq 83.38\%$ recall on channel 9
@@
Table I shows a comparison of the existing work with our approach. Most state-of-the-art solutions rely on the installation of fixed UWB radar systems. For example, Van Herbruggen et al. [11] introduce Accumulation Channel Impulse Response (ACIR) technique, which improves signal resolution for obstacle detection. Although this method increases the detection accuracy, it requires the collection of sufficient samples before it can be applied, limiting the maximum speed of robotic platforms. Sattiraju et al. [12] conclude that Random Forest and Extra Trees achieve high accuracy in obstacle detection. However, such learning-based methods are prone to overfitting the training data, which can reduce generalisation to new environments. Yun et al. [13] suppress noise from the received signal using singular value decomposition (SVD). Although their method achieves high accuracy, the required matrix computations are computationally intensive and may increase the latency and memory usage for real-time mobile robots. Mimouna et al. [14], [15] focus on obstacle detection, but lack characterisation of the obstacle (e.g. estimation of distance and width). Moreover, the use of complex deep learning models in [14] would lead to a higher execution time.
In general, learning-based methods are well suited for capturing complex signal characteristics and can achieve high detection accuracy when trained on sufficiently large and representative datasets. However, their performance strongly depends on the quality and diversity of the training data and may degrade when deployed in environments that differ from those seen during training. As a result, these methods are more prone to overfitting and reduced generalisation in unseen environments.
C. Smeenk et al. [16] present a signal processing chain for M-sequence UWB radars mounted on a mobile robot. Their approach combines background subtraction to remove the antenna crosstalk, with a constant false alarm rate (CFAR) method for object detection and distance estimation. However, the localisation is limited to reporting that an object is in front of the antenna and its distance, without providing its relative position in the environment. The method presented in this work overcomes all the above limitations by using clustering that is more adaptable to unseen environments. Moreover, our work also characterizes the position and width of detected obstacles by combining both distance estimation and AoA information.
@@
To the best of our knowledge, only four prior publications focus on robot-mounted UWB radar systems. [18] and [19] both focus more on Simultaneous Localisation and Mapping (SLAM) localisation aspects than UWB signal processing optimisations for obstacle characterisation. They are also limited by recognizing only vertical rods rather than reconstructing full obstacle representations from various material types. The proposed approach from our work can serve as a complementary method to [18] and [19], aiming to suppress noise and mitigate ghost detections while increasing the overall performance of their proposed UWB-SLAM algorithm. Unlike [19], which complements UWB-SLAM with localisation by employing AoA nodes for robot localisation and loop closure, our paper focuses on estimating the AoA of detected objects using the PDoA between antennas on the same chip. R. Zetik et al. [17] also focus on SLAM and demonstrate good localisation accuracy across different room shapes. However, their approach relies on proprietary antenna designs tailored for their experiments, which increases deployment cost.
@@
Finally, almost all prior work relies on custom UWB radar-only hardware devices with larger bandwidths and thus higher range resolution. Instead, our proposed approach is fully compliant with low-cost IEEE 802.15.4 standard off-the-shelf hardware. Using this setup, the same hardware can also localise the robot if fixed infrastructure is available, by reusing the transmitted wireless signal for multiple purposes simultaneously. This results in reduced energy consumption, reduced need for separate hardware, lower costs and weight, and more efficient use of the sparse spectrum. In this way, low-cost joint sensing, communication, and localisation for robotic platforms can be realised [20].
@@ IV Proposed approach
Figure 2: Overview of the proposed approach: Each capture contains the preamble CIR and two additional CIRs from both antennas on the same chip used for calculating the PDoA. A new measurement happens every 10.42 ms and contains the data of the environment at one timestamp. These are processed independently followed by filtering to remove noise and unwanted multipath components using peak properties, SNR-score, and PDoA. The distance to the object is estimated and the AoA is then calculated. Subsequently, multiple rows are combined to cluster reflections from the same object and map the environment.